The Shirt Works sells a large variety of tee shirts and sweatshirts. Steve Hooper, the owner, is thinking of expanding his sales by hiring high school students, on a commission basis, to sell sweatshirts bearing the name and mascot of the local high school.

       These sweatshirts would have to be ordered from the manufacturer six weeks in advance, and they could not be returned because of the unique printing required. The sweatshirts would cost Hooper $21.00 each with a minimum order of 360 sweatshirts. Any additional sweatshirts would have to be ordered in increments of 50.

       Since Hooper’s plan would not require any additional facilities, the only costs associated with the project would be the costs of the sweatshirts and the costs of the sales commissions. The selling price of the sweatshirts would be $42.00 each. Hooper would pay the students a commission of $4.00 for each shirt sold.


Required:
1.

To make the project worthwhile, Hooper would require a $16,150 profit for the first three months of the venture. What level of unit sales and dollar sales would be required to reach this target net operating income? (Round your intermediate calculations to 2 decimal places.)
Answer #1

1)

Unit sales to achieve target operating income =[Fixed cost+ target income ]/contribution per shirt

                       = [0+16150 ] / 17

                      = 16150 /17

                      = 950 shirts

Dollar sales to achieve target income = [Fixed cost+ target income ]/contribution margin ratio

                          = [0 +16150 ]/.4048

                           =$ 39896 (approx to 39900,difference is due to rounding off decimal places in contribution margin ratio )

Working:

Total variable cost= cost of shirt + sales commission

                     = 21 + 4

                      = $ 25

contribution per shirt= price-variable cost

               = 42 -25

               = $ 17

contribution margin ratio =contribution /price

                           = 17 /42

                          = .4048 or 40.48%